# Generalized Linear Bandits with Limited Adaptivity

### v4

@@ E.1 Convex Relaxation
The optimization problem in (28) is a non-convex optimization problem and therefore it is not clear what is the computational complexity of the problem. However, it is possible to substitute this optimization problem with a convex one, whose computational complexity can be better tractable. The process is similar to the one detailed in [Abeille et al., 2021, section 6]. Here we briefly outline the steps.
Let ${\mathcal{L}}_{t}(\theta)=\sum_{s=1}^{t}\ell(\theta,x_{s},r_{s})$ and $\breve{\theta}$ be defined as follows:
@@
$\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R(c/R)}$ ($\max_{i\in[s]}|\langle x_{s},\widetilde{\theta}-\theta^{*}\rangle|\leq c/R$)
@@
Let us define $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\coloneqq\sum_{s=1}^{t}\tilde{\alpha}(x_{s},\theta^{*},\theta)x_{s}x_{s}^{\intercal}$. Using the above fact, we obtain $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\succeq\frac{1}{2+c}{\mathbf{H}}^{*}$.
We now follow closely the proof outlined in Appendix B.3 of Abeille et al. [2021] with minor changes. By second-order Taylor’s expansion, for any $\theta\in\mathbb{R}^{d}$, we can write
${\mathcal{L}}_{t}(\theta)-{\mathcal{L}}_{t}(\theta^{*})-\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\theta-\theta^{*}\rangle$ $=\left\lVert\theta-\theta^{*}\right\rVert_{\tilde{{\mathbf{G}}}(\theta,\theta^{*})}^{2}$
$\geq\frac{1}{2+c}\left\lVert\theta-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}$
@@
Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$. By a similar Taylor expansion as above and some algebraic manipulations (see Appendix B.3 of Abeille et al. [2021]), we have, for $\theta^{*}$.
${\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})$ $\leq\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}(\theta^{*},\widehat{\theta})^{-1}}^{2}$
$\leq\frac{R}{\sqrt{\lambda}}\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}^{2}+\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$

### v5

@@ E.1 Convex Relaxation
The optimization problem in (32) is a non-convex optimization problem and therefore it is not clear what is the computational complexity of the problem. However, it is possible to substitute this optimization problem with a convex one, whose computational complexity can be better tractable. The process is similar to the one detailed in [2, section 6]. Here we briefly outline the steps.
Let ${\mathcal{L}}_{t}(\theta)=\sum_{s=1}^{t}\ell(\theta,x_{s},r_{s})+\lambda\left\lVert\theta^{*}\right\rVert^{2}$ and $\breve{\theta}$ be defined as follows:
@@
$\geq\frac{\dot{\mu}\left(\langle x_{s},\theta^{*}\rangle\right)}{2+R(c/R)}$ ($\max_{i\in[s]}|\langle x_{s},\breve{\theta}-\theta^{*}\rangle|\leq c/R$)
@@
Let us define $\tilde{{\mathbf{G}}}(\theta^{*},\theta)\coloneqq\sum_{s=1}^{t}\tilde{\alpha}(x_{s},\theta^{*},\theta)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$. Using the above fact, we obtain $\tilde{{\mathbf{G}}}(\theta^{*},\breve{\theta})\succeq\frac{1}{2+c}{\mathbf{H}}^{*}$.
We now follow closely the proof outlined in Appendix B.3 of [2] with minor changes. By second-order Taylor’s expansion, we can write
${\mathcal{L}}_{t}(\breve{\theta})-{\mathcal{L}}_{t}(\theta^{*})-\langle\nabla{\mathcal{L}}_{t}(\theta^{*}),\theta-\theta^{*}\rangle$ $=\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{\tilde{{\mathbf{G}}}(\breve{\theta},\theta^{*})}^{2}$
$\geq\frac{1}{2+c}\left\lVert\breve{\theta}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}}^{2}$
@@
Recall that $\widehat{\theta}$ is the unconstrained MLE, therefore $\nabla{\mathcal{L}}_{t}(\widehat{\theta})=\mathbf{0}$. By a similar Taylor expansion as above and some algebraic manipulations (Lemma 2 in Appendix B.4  [^12] The proof proceeds analogously, except that we retain the upper bound in terms of $\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$ instead of substituting it with $\gamma(\delta)$ at the end. and Appendix B.3 of [2] ), we have, for $\theta^{*}$.
${\mathcal{L}}_{t}(\theta^{*})-{\mathcal{L}}_{t}(\widehat{\theta})$ $\leq\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{G}}(\theta^{*},\widehat{\theta})^{-1}}^{2}$ (see Appendix B.3 of [2])
$\leq\frac{R}{\sqrt{\lambda}}\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}^{2}+\left\lVert g(\theta^{*})-g(\widehat{\theta})\right\rVert_{{\mathbf{H}}^{*-1}}$ (Lemma 2 in Appendix B.4 [2] )